OpenStudy (harrisonshark):

Select the correct product. (x 2 + 6x + 9)(3x - 1)

OpenStudy (harrisonshark):

its x^2

OpenStudy (harrisonshark):

@zaibali.qasmi

OpenStudy (harrisonshark):

3x^3 + 17x^2 + 21x - 9 3x^3 - 17x^2 - 21x - 9 3x^3+ 19x^2 + 27x + 9 3x^3+ 19x^2 + 9x - 9

OpenStudy (harrisonshark):

options ^

OpenStudy (anonymous):

(x 2 + 6x + 9)(3x - 1) =3x^3+18x^2+27x-x^2-6x-9 =3x^3+17x^2+21x-9 option A

OpenStudy (the_fizicx99):

FOIL is out, (x^2(3x -1)) + (6x(3x - 1)) + (9(3x - 1)) \(\ \sf \color{blue}{x^2}(\color{red}{3x - 1}) \implies (\color{blue}{x^2} \times\color{red}{3x}) + (\color{blue}{x^2} \times \color{red}{-1}) = 3x^3 + (-x^2) = 3x^3 - x^2 \)

OpenStudy (harrisonshark):

is he still right? @tHe_FiZiCx99

OpenStudy (the_fizicx99):

Option A is correct. However you shouldn't take answers that are given our directly. There's a high possibility that it could be wrong.